# Calculating the EFT likelihood via saddle-point expansion

### v1

@@ 2.2 Quantum field theory case
Substitute the field expansion into the path integral, we can get
$\mathcal{Z}$ $=\int\mathcal{D}\phi\,e^{-S_{E}[\phi]}$
$\approx\int_{C}\mathcal{D}\varphi\,e^{-S_{E}[\bar{\phi}]-\frac{1}{2}S^{\prime\prime}_{E}[\bar{\phi}]\varphi^{2}}$
$\sim e^{-S_{E}[\bar{\phi}]}\left({\rm det}S^{\prime\prime}[\bar{\phi}]\right)^{-1/2},$ (2.13)
where in the second line we have deformed the integral path into the steepest descent contours passing through the saddle points (labeled by $C$). This is because, in the case of quantum field theory, there is still a negative mode corresponding to the bounce solution (The exact bounce solution for some given potential can be found in [41, 42, 43, 44]). Since quantum field theory is just quantum mechanics, with the Hilbert space replaced by the infinite-dimensional Fock space, the application of the steepest descent contour method is not different in both cases. Therefore, we can adopt the same argument as in Sec. 2.1. In a general false vacuum decay theory, there are usually three types of saddle points of $S_{E}[\phi]$: false vacuum static (FV), bounce (B) and shot (S). The answer can thus be written as
$\mathcal{Z}$ $=\mathcal{J}_{\rm FV}+\mathcal{J}_{\rm B}+\mathcal{J}_{\rm S}$
$\sim\mathcal{I}_{\rm FV}+\mathcal{I}_{\rm B}+\mathcal{I}_{\rm S}\,.$ (2.14)
Note that in false vacuum decay theory, we should usually compare the contributions of the three components to the total path integral, with the bounce part $\mathcal{I}_{\rm B}$ ultimately being the dominant term. However, in this work we will focus only on the saddle-point expansion and the steepest descent contour method itself, and can omit this part without losing rigor.
@@ 3.1 Set up the path integral formula
We first consider the path integral approach to the matter and joint likelihood. After inheriting the generalization of the predecessors [21, 19, 47, 45, 46] (see also some functional methods in [48, 49, 50]), we write the partition function for the correlation functions of the matter field as
@@
Also at this point, $\cdots$ represents higher-order terms. A noteworthy point is that all the stochastic terms stem from the coarse-grained matter field, as these terms are required to counteract the UV dependence of the loop diagrams. Therefore, we need to regularize the integral over $\mathcal{D}\delta_{\rm in}$ by introducing a hard cutoff in the Fourier space. The relevant content has been illustrated in detail in the previous literature [51, 19], we will skip this discussion and follow the convention of [19].
A prominent advantage of the path-integral formulation is that it allows us to directly obtain the expression for the likelihood without making approximations, as compared to [20]. The likelihood is simply the functional Fourier transform of the partition function. If we define the functional Dirac function with a normalization factor $\mathcal{N}_{\delta^{(\infty)}}$,
@@
We can then derive the functional formula for the matter likelihood via Eq. (3.1)
@@
where we have written the current corresponding to $\delta$ as $X=iJ$. Similarly, we can obtain the expression of the joint likelihood
@@ 3.2 The saddle-point expansion
where we can read off the expression of $\mathcal{J}$ through Eq. (3.5), $\mathcal{J}=(i\delta,0)$, and the remaining matrix expressions can be extracted in the same manner. There are three things that should be emphasized about the above equation: First, in contrast to [19], the linear term of $\boldsymbol{\phi}$ has been incorporated into the action. Note that when performing the saddle-point expansion, the saddle point must be computed by considering all terms in the exponent, therefore under the former convention the saddle point of the partition function corresponding to the connected diagrams is required. Second, when calculating the decay rates, the first step is to transform to the Euclidean space. The equation of motion in Euclidean space admits an additional bounce solution compared to the Minkowski spacetime. This solution becomes the dominant contribution to the path integral and does not appear in the normalization constant [31]. However, in the present case, we compute the likelihood directly in the three-dimensional Euclidean space, so there is no need to consider the issue of the extra instanton solution. In the following calculations, we will also omit the normalization constant $\mathcal{N}_{\delta^{(\infty)}}$ which is independent of the fields, and recover it at the end. Another point worth noting is that, as seen from Eq. (3.5) and Eq. (3.6), the expression of the actions may not be real, since the imaginary unit $i$ appears in the functional Fourier transformation. Therefore, one may think that the result of the likelihood is not real. However, we reiterate here that the likelihood is real not only because it needs to correspond to the probability, but also, as we will demonstrate in the following example, if we utilize the gradient flow equation to solve for the steepest descent contour, all eigenvalues will also lead to a real likelihood.
Now that we have the consistent formula, we can compute this “path integral” as we did in Sec. 2. We split the field $\boldsymbol{\phi}$ into two parts $\boldsymbol{\phi}=\bar{\boldsymbol{\phi}}+\boldsymbol{\varphi}$, in which $\bar{\boldsymbol{\phi}}$ is the saddle point of the action, satisfying
@@
and $\boldsymbol{\varphi}$ is the “quantum fluctuation”. Then the likelihood can be approximated as
@@
If there are negative modes in the computation process, i.e. there are negative eigenvalues of $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$, then our calculation will also be ill-defined. At this point, we need to utilize the same argument as in quantum mechanics and quantum field theory, approximating the integral contour by the combination of the steepest descent contours connect different saddle points and terminate at the convergence regions in the $\boldsymbol{\phi}$ space. Thus we have
@@
where $C$ is one of the choices of the steepest descent contour that is homologous to the initial integral contour we have chosen. It can be seen that the situation here is analogous to that in quantum field theory when calculating the decay rate with multiple fields. So after proceeding with this, we can apply the same discussion as in quantum field theory.
@@
Another naive problem is whether it is justified to use the saddle-point expansion in the path-integral approach of the LSS. In quantum field theory, if we recover $\hbar$, then the field decomposition should be $\phi=\bar{\phi}+\hbar^{1/2}\varphi$, such that the contribution near the saddle point is of order $\mathcal{O}(\hbar^{-1})$ (i.e. proportional to $\hbar^{-1}\varphi^{2}$), and all higher-order terms are suppressed by powers of $\hbar$. However now the path-integral approach of LSS is not a quantum theory, so there is no $\hbar$ in the expressions for $\boldsymbol{\phi}$ or $\boldsymbol{\phi}_{g}$, nor is there a similar perturbative parameter in this theory which can be used to draw an analogy. In fact, one compelling piece of evidence supporting the validity of the saddle-point approximation in this context is its application in [19], where it is employed to compute the likelihoods, yielding results that are not only consistent with those derived from the traditional EFT method [20], but also provide some additional higher-order terms. Therefore, it is reasonable to conclude that the contributions from the saddle points will dominate the entire path integral. Furthermore, the contributions near the saddle points can then be appropriately interpreted as the second-order corrections to the likelihoods. We acknowledge that the functional approach based on the path integral provides more accurate results, and our method will further advance this framework. In the subsequent section, we will illustrate that the application of the saddle-point expansion introduces an additional factor in the likelihood result than before, which aligns with our expectations.
4 Example: the EFT likelihood with Gaussian distribution
@@
Upon establishing the appropriate theoretical framework, it can be applied to specific models. In this section, we focus on the model presented in [19], which involves the computation of the conditional likelihood with Gaussian distribution. We then demonstrate the application of the steepest descent contour through concrete examples.
@@
As outlined in Sec. 3.1, there are three stochastic terms for matter and galaxy, $P_{\epsilon_{g}}$, $P_{\epsilon_{g}\epsilon_{m}}$, $P_{\epsilon_{m}}$. In the case where only $P_{\epsilon_{g}}$ exists, the path integral can be evaluated exactly, yielding rigorous results. This example has been previously investigated in [20, 19]. Hence, we consider the more intricate case where both $P_{\epsilon_{g}}$ and $P_{\epsilon_{g}\epsilon_{m}}$ are non-vanishing, and demonstrate that the saddle-point expansion yields identical conditional likelihood.
@@
We can read off the formula of $\mathcal{M}^{ab}$ and $\mathcal{M}_{g}^{ab}$ from the expression of the likelihood, i.e. Eq. (3.5) and Eq. (3.6)
@@
At zeroth order, we can take $K_{1}=1$ and $K_{g,1}=b_{1}$. We briefly review how to calculate the contribution of the saddle points, with the detailed results provided in Appendix A, and here we focus on the issues related to the negative modes.
@@
Note that due to the presence of the Dirac delta function in the expressions for $\mathcal{M}$ and $\mathcal{M}_{g}$, we need not consider the case when $\boldsymbol{k}\neq\boldsymbol{k}^{\prime}$. The expressions of the solutions of these two equations and $-S[\bar{\boldsymbol{\phi}}]$, $-S_{g}[\bar{\boldsymbol{\phi}}_{g}]$ can be found in Appendix A. We now consider the additional terms arising from the saddle-point expansion. For both the matter and galaxy case, $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$ and $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$ are simply the matrices corresponding to the second-order terms in the fields, $\mathcal{M}^{ab}$ and $\mathcal{M}^{ab}_{g}$. For the matter likelihood, the eigenvalues of $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$ can be simply calculated,
@@
$\mathcal{P}[\delta]\approx e^{-S[\bar{\boldsymbol{\phi}}]}({\rm det}S^{\prime\prime}[\bar{\boldsymbol{\phi}}])^{-1/2}=\sqrt{\pi}\,(\prod_{i}m_{i})^{-1/2}\,{\rm e}^{-S[\bar{\boldsymbol{\phi}}]}=\frac{\sqrt{\pi}}{2}D_{1}\,e^{-S[\bar{\boldsymbol{\phi}}]}.$ (4.8)
It can be seen that after incorporating the second-order expansion of the saddle points, the resulting expression is equivalent to the original one with the addition of an extra factor, as anticipated. Analogously, the calculation of joint likelihood should be obtained in the same manner, however, we will encounter certain intractable problems now: For a general three-dimensional matrix $\mathcal{M}_{g}$, there are no simple analytical expressions for its eigenvalues. Even when we obtain them through complicated derivations, the solutions are usually not positive, and even not real because $\mathcal{M}_{g}$ is neither Hermitian nor anti-Hermitian. At this time, there are no shortcuts, and we need to follow the previous discussions regarding the negative modes. We need to complexify the field $\boldsymbol{\phi}_{g}$ and then seek the steepest descent contour that is homologous to the original contour. At the same time, there is no zero mode corresponding to $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$, so we do not need to be concerned about the issues related to the collective coordinates.
We will present a general method for solving this type of problem, and the following discussion is derived by analogy with [35]. We start from the complexified $\boldsymbol{\phi}$ space. We introduce the real parameter $u$ to label different paths passing through the saddle points in the complex plane, and impose the boundary condition $\boldsymbol{\phi}_{g}(\boldsymbol{k},u\rightarrow\infty)=\bar{\boldsymbol{\phi}}_{g}$, i.e. under this parametrization, in the limit of $u\rightarrow\infty$, we recover the saddle points. Since our goal is to identify the steepest descent contours, the holomorphic function we consider should be the negative value of the action $\mathcal{I}[\boldsymbol{\phi}_{g}]=-S_{g}[\boldsymbol{\phi}_{g}]$, and then define the Morse function $f[\boldsymbol{\phi}_{g}]={\rm Re}(\mathcal{I}[\boldsymbol{\phi}_{g}])$. Thus, for the saddle point $\boldsymbol{\phi}_{g}$, the steepest descent path is given by the gradient flow equation [34]
@@
where the overline represents the complex conjugate. We can easily check that
@@
thus the operator on the left-hand side is now Hermitian, and the eigenvalues $m_{n}$ should be real. However, this equation remains intractable for an exact solution, and the steps for obtaining an approximate solution are provided in Appendix B. In this case, we can also impose the orthonormalization to the eigenstates
@@
On the other hand, for the equation on the RHS of Eq. (4.14), the solution is expected to take the form $g_{n}(u)\sim a_{n}\,{\rm exp}\,(m_{n}u)$, where $a_{n}\in\mathbb{R}$. Using the boundary condition $g(u\rightarrow\infty)=0$, it follows that $m_{n}>0$, i.e. the eigenvalues are all positive and real.
@@
$\mathcal{P}[\delta_{g},\delta]$ $\approx\int_{C}\mathcal{D}\boldsymbol{\phi}_{g}\,e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]-\frac{1}{2}\boldsymbol{\varphi}_{g}S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]\boldsymbol{\varphi}_{g}}$
$=\int_{C}\mathcal{D}\boldsymbol{\varphi}_{g}\,e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]-\boldsymbol{\varphi}_{g}M_{g}\boldsymbol{\varphi}_{g}}$
$=\int\mathcal{D}\boldsymbol{\varphi}_{g}\,e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]-\sum_{n}m_{n}|\boldsymbol{\varphi}_{g}|^{2}}$
$=e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]}\frac{\pi}{\sum_{n}m_{n}}\,$
$\approx e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]}\frac{\pi}{P_{\epsilon_{g}}-P_{\rm in}^{-1}},$ (4.18)
where from the second to the third line we have employed the eigenequation of $\mathcal{M}_{g}$, and in the subsequent line the Gaussian integral formula for complex variables has been applied. In the final step, we took advantage of the expression of the sum of $m_{n}$, i.e. Eq. (B.9).

### v2

@@ 2.2 Quantum field theory case
In the case of quantum field theory, there are still three types of saddle points of $S_{E}[\phi]$: false vacuum static (FV), bounce (B) and shot (S). Substitute the field expansion into the path integral, we can get
$\mathcal{Z}=\int\mathcal{D}\phi\,e^{-S_{E}[\phi]}$ $\approx\sum_{n=FV,B,S}\int_{C_{n}}\mathcal{D}\varphi_{n}\,e^{-S_{E}[\bar{\phi}_{n}]-\frac{1}{2}S^{\prime\prime}_{E}[\bar{\phi}_{n}]\varphi^{2}_{n}}\,,$ (2.14)
where in the second step we have deformed the integral path into the steepest descent contours passing through the saddle points (labeled by $C_{n}$). This is because, in the case of quantum field theory, there is still a negative mode corresponding to the bounce solution (The exact bounce solution for some given potential can be found in Refs. [42, 43, 44, 45]). Since quantum field theory is just quantum mechanics, with the Hilbert space replaced by the infinite-dimensional Fock space, the application of the steepest descent contour method is not different in both cases. Therefore, we can adopt the same argument as in Sec. 2.1. Thus, the answer can be written as
$\mathcal{Z}=\mathcal{J}_{FV}+\mathcal{J}_{B}+\mathcal{J}_{S}\sim\mathcal{I}_{FV}+\mathcal{I}_{B}+\mathcal{I}_{S}\,.$ (2.15)
Note that in false vacuum decay theory, we should usually compare the contributions of the three components to the total path integral, with the bounce part $\mathcal{I}_{\rm B}$ ultimately being the dominant term. A detailed discussion of this process can be found in [36, 46]. Since the application of the saddle-point expansion in this work does not involve multiple saddle points, we can safely omit this part without compromising rigor.
@@ 3.1 Set up the path integral formula
We first consider the path integral approach to the matter and joint likelihoods. After inheriting the generalization of the predecessors [21, 19, 49, 47, 48] (see also some functional methods in Refs. [50, 51, 52]), we write the partition function for the correlation functions of the matter field as
@@
Also at this point, $\cdots$ represents higher-order terms. A noteworthy point is that all the stochastic terms stem from the coarse-grained matter field, as these terms are required to counteract the UV dependence of the loop diagrams. Therefore, we need to regularize the integral over $\mathcal{D}\delta_{\rm in}$ by introducing a hard cutoff in the Fourier space. The relevant content has been illustrated in detail in the previous literatures [53, 19], we will skip this discussion and follow the convention of [19].
A key advantage of the path-integral formulation is that it allows us to exploit all available information. In [19], the authors calculated the conditional likelihood using this approach and demonstrated that, under the same assumptions — such as the Gaussian initial condition and two nonvanishing stochasticities $P_{\epsilon_{g}}$ and $P_{\epsilon_{g}\epsilon_{m}}$ — the path-integral approach not only reproduces the results of [20], but also recovers certain terms that were omitted in that work.
The origin of these additional terms can be attributed to a drawback of the standard perturbation theory (SPT): when computing the $n$-point correlation functions, correlators of order higher than $n$ are involved. This can be seen, for example, in the equations referred to as the BBGKY hierarchy in [54]. At the same time, from the continuity equation for the probability density in [50] (where this approach is also called the time-sliced perturbation theory) and the renormalization group equations in Refs. [49, 47, 48], it is evident that in this path-integral formulation, the $n$-th order evolution equation will not contain terms higher than $n$. Therefore, we conclude that these extra terms are indeed missing in [20], and the functional method provides more precise results.
In this approach, the likelihood is simply the functional Fourier transform of the partition function. If we define the functional Dirac function with a normalization factor $\mathcal{N}_{\delta^{(\infty)}}$,
@@
We can then derive the functional formula for the matter likelihood using Eq. (3.1)
@@
where we have written the current corresponding to $\delta$ as $X=iJ$. Similarly, we can obtain the expression for the joint likelihood
@@ 3.2 The saddle-point expansion
where we can read off the expression for $\mathcal{J}$ through Eq. (3.5), $\mathcal{J}=(i\delta,0)$, and the remaining matrix expressions can be extracted in the same manner. There are three things that should be emphasized about the above equation: First, in contrast to [19], the linear term of $\boldsymbol{\phi}$ has been incorporated into the action. Note that when performing the saddle-point expansion, the saddle point must be computed by considering all terms in the exponent, therefore under the former convention the saddle point of the partition function corresponding to the connected diagrams is required. Second, when calculating the decay rates, the first step is to transform to Euclidean space. The equation of motion in Euclidean space admits an additional bounce solution compared to the Minkowski spacetime. This solution becomes the dominant contribution to the path integral and does not appear in the normalization constant [31]. However, in the present case, we compute the likelihood directly in the three-dimensional Euclidean space, so there is no need to consider the issue of the extra instanton solution. In the following calculations, we will also omit the normalization factor $\mathcal{N}_{\delta^{(\infty)}}$ in Eq. (3.4) which is independent of the fields, and recover it at the end. Another point worth noting is that, as seen from Eq. (3.5) and Eq. (3.6), the expression for the actions may not be real, since the imaginary unit $i$ appears in the functional Fourier transformation. Therefore, one may think that the results of the likelihoods may not be real. However, we reiterate here that the likelihoods are real not only because they need to correspond to the probability, but also, as we will demonstrate in the following example in Sec. 4, if we use the gradient flow equation to solve for the steepest descent contours, all eigenvalues will also lead to a real conditional likelihood.
Now that we have the consistent formula, we can compute this “path integral” as we did in Sec. 2. We decompose the field $\boldsymbol{\phi}$ into two parts $\boldsymbol{\phi}=\bar{\boldsymbol{\phi}}+\boldsymbol{\varphi}$, in which $\bar{\boldsymbol{\phi}}$ is the saddle point of the action, satisfying
@@
and $\boldsymbol{\varphi}$ is the “quantum fluctuation”. Then the matter likelihood can be approximated as
@@
If there are negative modes in the computation process, i.e. there are negative eigenvalues of $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$, then our calculation will also be ill-defined. At this point, we need to use the same argument as in quantum mechanics and quantum field theory, approximating the integral contour by the combination of the steepest descent contours connect different saddle points and terminate at the convergent regions in the $\boldsymbol{\phi}$ space. Thus we have
@@
where $C$ is one of the choices of the combination of the steepest descent contours that is homologous to the original integral contour we have chosen. It can be seen that the situation here is analogous to that in quantum field theory when calculating the decay rate with multiple fields [55]. So after proceeding with this, we can apply the same discussion as in quantum field theory.
@@
Another naive problem is whether it is justified to use the saddle-point expansion in the path-integral approach to the LSS. In quantum field theory, if we recover $\hbar$, then the field decomposition should be $\phi=\bar{\phi}+\hbar^{1/2}\varphi$, such that the contribution near the saddle point is of order $\mathcal{O}(\hbar)$ (i.e. proportional to $\hbar\,\varphi^{2}$), and all higher-order terms are suppressed by powers of $\hbar$. However now the path-integral approach to the LSS is not a quantum theory, so there is no $\hbar$ in the expression for the decomposition of $\boldsymbol{\phi}$ or $\boldsymbol{\phi}_{g}$, nor is there a similar perturbative parameter in this theory which can be used to draw an analogy.
We hereby offer a justification for using saddle points to approximate the full path integral. The rationale can be traced to the method of stationary phase (also known as the method of steepest descent) in quantum field theory (see, for example, Sec. 14.2 of [56]). Suppose we consider a path integral of the following form
$\int\mathcal{D}\Phi(\boldsymbol{x},t)\,e^{-\frac{1}{\hbar}S[\Phi]}.$ (3.17)
In the classical limit $\hbar\rightarrow 0$, this integral is dominated by the values of $\Phi$ for which $S[\Phi]$ has an extremum, i.e. $\delta S[\Phi]=0$, because these values correspond to the regions where the action oscillates at least. This is why we can decompose the field as $\Phi=\bar{\Phi}+\mathcal{O}(\hbar)\,\phi$, where $\bar{\Phi}$ is the stationary point of the action. We expect the same phenomenon to occur in our case. This can be achieved by performing a rescaling analogous to the one in [19], taking $\boldsymbol{k}=b\boldsymbol{k}^{\prime}$ and setting the “external currents” $\delta$ and $\delta_{g}$ to zero. Therefore, in the limit $b\rightarrow 0^{+}$, the rescaling of different operators will reflect their relevance (or irrelevance) to the large-scale universe. In this context, we employ $b$ [^4] Note that here $b$ is not a perturbative parameter but a rescaling parameter. as a parameter analogous to $\hbar$ in quantum mechanics. Under this rescaling, the integration measure ${\rm d}^{3}k$ changes to $b^{3}{\rm d}^{3}k^{\prime}$, and the power spectrum $P_{\rm in}$ scales as $P_{\rm in}(k)=b^{n_{\delta}}P_{\rm in}(k^{\prime})$, where $n_{\delta}$ is the order of momentum in the initial power spectrum. Then, the terms in the action that are quadratic in $X_{g}$ and $\delta_{\rm in}$ remain invariant if we redefine
$X_{g}(b\boldsymbol{k}^{\prime})$ $=b^{-\frac{3}{2}}X_{g}^{\prime}(\boldsymbol{k^{\prime}})\,,$ (3.18)
$\delta_{\rm in}(b\boldsymbol{k}^{\prime})$ $=b^{\frac{n_{\delta}-3}{2}}\delta_{\rm in}(\boldsymbol{k}^{\prime})\,.$ (3.19)
For $n_{\delta}=-1.7$, it can be observed that the rescaling rate of $\delta_{\rm in}$ is faster than $X_{g}$. In this case, if we only consider the lowest-order nonvanishing terms in the action Eq. (3.15), then the rescaling behavior of the action is
$S_{g}[\boldsymbol{\phi}_{g}]\rightarrow b^{-3}S_{g}[\boldsymbol{\phi}_{g}]\,,$ (3.20)
and the path-integral should to be rescaled as
$\int\mathcal{D}\boldsymbol{\phi}_{g}\,{\rm}e^{-S_{g}[\boldsymbol{\phi}_{g}]}\rightarrow\int\mathcal{D}\boldsymbol{\phi}_{g}\,{\rm}e^{-\frac{1}{b^{3}}S_{g}[\boldsymbol{\phi}_{g}]}\,.$ (3.21)
We can now draw an analogy with the situation in quantum field theory: at large scales ($b\rightarrow 0^{+}$), the saddle-point contribution to the action will dominate the path integral, and other terms are suppressed. Thus, we have demonstrated that it is convincing to approximate the path integral using the saddle points. Then, the field expansion should be
$\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}+\mathcal{O}(b)\,\boldsymbol{\varphi}_{g}\,.$ (3.22)
Furthermore, the contributions near the saddle points can then be appropriately interpreted as the second-order corrections to the likelihoods. We acknowledge that the functional approach based on the path integral provides more accurate results, and our method will further advance this framework. In the subsequent section, we will illustrate that the application of the saddle-point expansion introduces an additional factor in the likelihood result than before, which aligns with our expectations.
4 Example: the EFT likelihood with a Gaussian distribution
@@
Upon establishing the appropriate theoretical framework, it can be applied to specific models. In this section, we focus on the model presented in [19], which involves the computation of the conditional likelihood with a Gaussian distribution. We then demonstrate the application of the steepest descent contour method through concrete examples.
@@
As outlined in Sec. 3.1, there are three stochastic terms for matter and galaxy, $P_{\epsilon_{g}}$, $P_{\epsilon_{g}\epsilon_{m}}$, $P_{\epsilon_{m}}$. In the case where only $P_{\epsilon_{g}}$ exists, the likelihoods can be evaluated exactly, yielding rigorous results. This example has been previously investigated in [20, 19]. Hence, we consider the more intricate case where both $P_{\epsilon_{g}}$ and $P_{\epsilon_{g}\epsilon_{m}}$ are non-vanishing, and demonstrate that the saddle-point expansion yields identical conditional likelihood.
@@
We can read off the formula of $\mathcal{M}^{ab}$ and $\mathcal{M}_{g}^{ab}$ from the expression for the likelihoods, i.e. Eq. (3.5) and Eq. (3.6)
@@
At zeroth order, we can take $K_{1}=1$ and $K_{g,1}=b_{1}$. We briefly review how to calculate the contribution of the saddle points, with the detailed results provided in Appendix B, and here we focus on the issues related to the negative modes.
@@
Note that due to the presence of the Dirac delta function in the expressions for $\mathcal{M}$ and $\mathcal{M}_{g}$, we need not consider the case when $\boldsymbol{k}\neq\boldsymbol{k}^{\prime}$. The expressions for the solutions of these two equations and $-S[\bar{\boldsymbol{\phi}}]$, $-S_{g}[\bar{\boldsymbol{\phi}}_{g}]$ can be found in Appendix B. We now consider the additional terms arising from the saddle-point expansion. For both the matter and galaxy case, $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$ and $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$ are simply the matrices corresponding to the second-order terms in the fields, $\mathcal{M}^{ab}$ and $\mathcal{M}^{ab}_{g}$. For the matter likelihood, the eigenvalues of $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$ can be simply calculated,
@@
$\mathcal{P}[\delta]\approx e^{-S[\bar{\boldsymbol{\phi}}]}({\rm det}S^{\prime\prime}[\bar{\boldsymbol{\phi}}])^{-1/2}=\sqrt{\pi}\,(\prod_{i}m_{i})^{-1/2}\,{\rm e}^{-S[\bar{\boldsymbol{\phi}}]}=\frac{\sqrt{\pi}}{D_{1}}\,e^{-S[\bar{\boldsymbol{\phi}}]}.$ (4.8)
It can be seen that after incorporating the second-order expansion of the saddle points, the resulting expression is equivalent to the original one with the addition of an extra factor, as anticipated. Analogously, the calculation of joint likelihood should be obtained in the same manner, however, we will encounter certain intractable problems now: Given a general three-dimensional matrix $\mathcal{M}_{g}$, there are no simple analytical expressions for its eigenvalues. Even when we obtain them through complicated derivations, the solutions are usually not positive, and even not real because $\mathcal{M}_{g}$ is neither Hermitian nor anti-Hermitian. At this time, there are no shortcuts, and we need to follow the previous discussions regarding the negative modes. We need to complexify the field $\boldsymbol{\phi}_{g}$ and then seek the approximate contour that is homologous to the original one through the gradient flow equation. At the same time, there is no zero mode corresponding to $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$, so we do not need to be concerned about the issues related to the collective coordinates.
We will present a general method for solving this type of problem, and the following discussion is derived by analogy with [35]. We start from the complexified $\boldsymbol{\phi}$ space. We introduce the real parameter $u$ to label different paths passing through the saddle points in the complex plane, and impose the boundary condition $\boldsymbol{\phi}_{g}(\boldsymbol{k},u\rightarrow-\infty)=\bar{\boldsymbol{\phi}}_{g}$, i.e. under this parametrization, in the limit of $u\rightarrow-\infty$, we recover the saddle points. Since our goal is to identify the steepest descent contours, the holomorphic function we consider should be the negative value of the action $\mathcal{I}[\boldsymbol{\phi}_{g}]=-S_{g}[\boldsymbol{\phi}_{g}]$, and then define the Morse function $f[\boldsymbol{\phi}_{g}]={\rm Re}(\mathcal{I}[\boldsymbol{\phi}_{g}])$. Thus, for the saddle point $\bar{\boldsymbol{\phi}}_{g}$, the steepest descent path is given by the gradient flow equation [34]
@@
where $u\in\mathbb{R}$ and the boundary condition is $\boldsymbol{\phi}_{g}(\boldsymbol{k},u\rightarrow-\infty)=\bar{\boldsymbol{\phi}}_{g}$. The overline represents the complex conjugate. We can easily check that
@@
thus the operator on the left-hand side is now Hermitian, and the eigenvalues $m_{n}$ should be real. However, this equation remains intractable for an exact solution, and the steps for obtaining an approximate solution are provided in Appendix C. In this case, we can also impose the orthonormalization to the eigenstates
@@
On the other hand, for the equation on the RHS of Eq. (4.14), the solution is expected to take the form $g_{n}(u)\sim a_{n}\,{\rm exp}\,(m_{n}u)$, where $a_{n}\in\mathbb{R}$. Using the boundary condition $g(u\rightarrow-\infty)=0$, it follows that $m_{n}>0$, i.e. the eigenvalues are all positive and real.
@@
$\mathcal{P}[\delta_{g},\delta]$ $\approx\int_{C}\mathcal{D}\boldsymbol{\phi}_{g}\,e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]-\frac{1}{2}\boldsymbol{\varphi}_{g}S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]\boldsymbol{\varphi}_{g}}=\int_{C}\mathcal{D}\boldsymbol{\varphi}_{g}\,e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]-\boldsymbol{\varphi}_{g}M_{g}\boldsymbol{\varphi}_{g}}$
$=\int\mathcal{D}\boldsymbol{\varphi}_{g}\,e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]-\sum_{n}m_{n}|\boldsymbol{\varphi}_{g}|^{2}}=e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]}\frac{\pi}{\sum_{n}m_{n}}\approx e^{-S_{g}[\bar{\boldsymbol{\phi}}_{g}]}\frac{\pi}{P_{\epsilon_{g}}-P_{\rm in}^{-1}},$ (4.18)
where in the first line we have employed the eigenequation of $\mathcal{M}_{g}$, and in the subsequent line the Gaussian integral formula for complex variables has been applied. We reiterate that the integral computed in the above expression corresponds to the sum of $\mathcal{J}_{i}$ rather than $\mathcal{I}_{i}$ as discussed in Sec. 2, as the integral contour consists of all trajectories satisfying the gradient flow equation across the entire complex plane, i.e. the combination of different Lefschetz thimbles. In the final step, we took advantage of the expression for the sum of $m_{n}$, i.e. Eq. (C.9).